When you have an equation such as x + 1/2 = -3/2 how do you know which number to subtract or add from the left side and vice versa on the right side? Is there an easy way to determine it? 
Answer by edjones(8007)

x + 1/2 = -3/2
x=-4/2 subtract 1/2 from each side because you want the x to be alone on the left hand side.
x=-2
.
Ed